# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 6 Data Analysis
Unit 1 0.008 0.073 (0.0002, 0.177) 0.041 (0.0002, 0.157)
Unit 2 0.010 0.066 (0.0001, 0.173) 0.035 (0.0001, 0.143)
Unit 3 0.001 0.061 (0.0004, 0.164) 0.020 (0.0001, 0.086)
Unit 4 0.003 0.059 (0.0003, 0.161) 0.025 (0.0001, 0.107)
Unit 5 0.003 0.062 (0.0004, 0.169) 0.016 (0.0001, 0.066)
Unit 6 0.006 0.061 (0.0004, 0.165) 0.019 (0.0001, 0.078)
Unit 7 0.006 0.062 (0.0004, 0.172) 0.055 (0.0002, 0.168)
Unit 8 0.009 0.062 (0.0004, 0.167) 0.033 (0.0001, 0.132)
Unit 9 0.003 0.056 (0.0002, 0.163) 0.015 (0.0001, 0.063)
Unit 10 0.005 0.058 (0.0003, 0.159) 0.014 (0.0001, 0.058)
Unit 11 0.008 0.058 (0.0003, 0.159) 0.021 (0.0001, 0.087)
Unit 12 0.008 0.059 (0.0003, 0.168) 0.025 (0.0001, 0.101)
Unit 13 0.009 0.055 (0.0003, 0.158) 0.059 (0.0004, 0.189)
Unit 14 0.013 0.052 (0.0002, 0.157) 0.027 (0.0001, 0.114)
Unit 15 0.018 0.055 (0.0003, 0.161) 0.035 (0.0001, 0.144)
Unit 16 0.024 0.056 (0.0003, 0.165) 0.060 (0.0003, 0.223)
Unit 17 0.026 0.056 (0.0002, 0.167) 0.064 (0.0004, 0.211)
Unit 18 0.031 0.057 (0.0003, 0.170) 0.066 (0.0003, 0.211)
Unit 19 0.040 0.058 (0.0002, 0.168) 0.070 (0.0004, 0.221)
Unit 20 0.047 0.057 (0.0003, 0.169) 0.072 (0.0003, 0.228)
Unit 21 0.050 0.059 (0.0003, 0.180) 0.072 (0.0004, 0.232)
@@
We derive residual lifetime distribution corresponding to the last time point at which degrading unit has not crossed the threshold value. Predictive intervals are constructed using the 95% highest probability density region of the generated samples using TMCMC method with $p=\frac{1}{2}$. Table 6 presents residual lifetime estimates where the lower limit of the predictive interval is rounded off to four decimal places and rest of the values are rounded off to three decimal places along with the true values for each of 21 units. Histograms of the generated residual lifetimes of first three units along with the trace plot and autocorrelation plot for the generated samples for semi-parametric and parametric method is presented in Figure 8 and 9 respectively. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit. It is observed that the proposed semi-parametric method is performing better than the parametric method in almost all the cases in terms of predicting residual lifetime of the individual units. It is also evident that for most of the cases the produced predictive intervals for both the methods contain true value. Comparison of prediction error for the both parametric and semi-parametric method is provided in Appendix D.
@@ 7 Conclusion
The objective of this article is to predict residual lifetime of units in a heterogeneous situation. To deal with heterogeneity we proposed a Bayesian semi-parametric degradation model. In the first part we assumed a general path model where the random effects are modeled using the Dirichlet process mixture of normal distributions. Model hierarchy is represented according to Ishwaran & Zarrerpur [31]. Gibbs sampling is used to draw samples from the posterior distribution of the model parameters. In the second part of the problem we developed the residual lifetime distribution of each unit which depends on the samples drawn from posterior distribution. Finally we simulate samples from this distribution using MCMC technique and considered the sample median as the predicted residual lifetime of each unit. For implementation of the proposed method we use R programming language. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method with respect to the error criteria discussed earlier in this paper. In simulation study it is found that, in some of the cases our proposed method is performing well compared to the parametric method whereas there are some cases where parametric method performs better than the proposed method. Finally, we applied the model to Fatigue Crack-Size dataset to evaluate the performance of our model. In this case, it is found that the prediction is more accurate for the proposed semi-parametric method compared to the parametric method for most of the units. The proposed Bayesian semi-parametric degradation model may be used for similar problems like in medical diagnosis contexts as well.
Appendix A Histogram of Residual lifetime, Monitoring MCMC convergence for Bayesian semi-parametric method
A.1 Case 1
@@
Figure 10: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples for unit 4-10 by semi-parametric method. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
A.2 Case 2
Figure 11: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 12: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 11-18. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 13: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 19-26. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 14: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 27-30. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
A.3 Fatigue - Crack Size Data
Figure 15: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 4-5. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 16: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 6-13. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Figure 17: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by semi-parametric method for unit 14-21. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Appendix B Histogram of Residual lifetime, Monitoring MCMC convergence for Bayesian parametric method
B.1 Case 1
@@
Figure 18: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
B.2 Case 2
@@
Figure 19: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-10. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
@@
Figure 20: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 11-18. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
@@
Figure 21: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 19-26. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
@@
Figure 22: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 27-30. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
B.3 Fatigue - Crack Size Data
@@
Figure 23: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 4-5. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
@@
Figure 24: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 6-13. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
@@
Figure 25: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot produced based on generated samples by parametric method for unit 14-21. Histograms and autocorrelation plots are constructed based on the samples at lag size 10 for each unit.
Appendix C Monitoring MCMC convergence of model parameters
C.1 Case 1
@@
Figure 26: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,..,8$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
Figure 27: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{9},\beta_{10},\alpha,\sigma_{\epsilon}^{2}$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.
C.1.1 Case 2
@@
Figure 28: For each row, first two plots from the left side is trace plot and autocorrelation plot produced by semi-parametric method and last two plots denotes trace plot and autocorrelation plot produced by parametric method based on the generated samples for $\beta_{i}$, $i=1,..,3$ respectively. Autocorrelation plots are constructed based on the samples at lag size 50.

### v2

@@ 6.4 Case 4
1 0.189 0.475 (0.073, 0.842) 0.478 (0.090, 0.846)
2 0.277 0.668 (0.206, 1.177) 0.667 (0.117, 1.084)
3 1.653 2.272 (1.187, 3.689) 2.285 (1.228, 3.692)
4 0.632 1.033 (0.350, 1.662) 1.033 (0.390, 1.694)
5 0.905 1.584 (0.797, 2.523) 1.541 (0.734, 2.466)
6 0.017 0.206 (0.001, 0.469) 0.205 (0.001, 0.458)
7 0.928 1.761 (0.893, 2.838) 1.773 (0.971, 2.894)
8 0.976 1.652 (0.809, 2.676) 1.652 (0.772, 2.670)
9 0.027 0.239 (0.001, 0.501) 0.235 (0.001, 0.499)
10 0.027 0.221 (0.001, 0.461) 0.222 (0.008, 0.478)
Table 22: KS distance between true and approximated residual lifetime distribution for Bayesian parametric method and proposed Bayesian semi-parametric method
. Unit $m=11$ $m=31$ Parametric Semi-parametric Parametric Semi-parametric 1 0.151 0.171 0.222 0.220 2 0.421 0.411 0.205 0.204 3 0.571 0.552 0.241 0.235 4 0.276 0.263 0.175 0.170 5 0.198 0.173 0.190 0.182 6 0.321 0.354 0.068 0.059 7 0.454 0.458 0.125 0.136 8 0.626 0.639 0.136 0.131 9 0.213 0.213 0.143 0.139 10 0.064 0.086 0.070 0.062
Table 23: Prediction accuracy for Parametric and Semi-Parametric method
Scenario $M_{1}$ $M_{2}$
Parametric Semi-parametric Parametric Semi-parametric
RMSE MAE RMSE MAE RMSE MAE RMSE MAE
$n=10,m=11$ 0.637 0.688 0.632 0.685 0.605 0.674 0.611 0.674
$n=10,m=31$ 0.501 0.669 0.499 0.668 0.486 0.657 0.492 0.660
$n=30,m=11$ 0.792 0.780 0.829 0.791 0.782 0.774 0.822 0.785
$n=30,m=31$ 0.720 0.742 0.725 0.742 0.726 0.740 0.738 0.743
$n=50,m=11$ 0.617 0.685 0.659 0.706 0.611 0.681 0.651 0.701
$n=50,m=31$ 0.620 0.722 0.621 0.721 0.625 0.723 0.626 0.722
@@
We present predictive intervals of residual lifetime for the individual units in Table 21. The estimates for residual lifetime are rounded off to three decimal places for each unit. The histograms of the generated residual lifetimes for units 1-3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method are provided in Figures 9 and 10 respectively in the Appendix A to this article. For most of the units, the produced predictive intervals for both parametric and semi-parametric methods contain the true value.
Kolmogorov-Smirnov (KS) statistic obtained for $m=11,31$ and $n=10$ for this case is given in Table 22. For most of the slowly degrading units for example unit 3, 8 value of KS statistic is less for semi-parametric method compared to parametric method for $m=11,31$. The semi-parametric model is able to accurately estimate the random effects associated with these units, due to the its multi-modality characteristics. For the moderate degrading units for example, unit 4 semi-parametric methods also produces lesser KS value. On the other side, for fastest degrading units the results are very similar. To asses the overall predictive performance of both the models, RMSE and MAE are calculated for each combination of $n$ and $m$. It is evident from Table 23 that, as the number of observations per unit increases for $n=10,30,50$, both RMSE and MAE decreases. It is found that the produced errors seems to similar with respect to both model under $M_{1}$ and $M_{2}$ as the number of units and observations per unit increases.
6.5 Case 5
Now, to asses the impact of proposed semi-parametric method, we consider a 5-component mixture of normal distribution for random effect. The degradation paths for the units generated for this case is given by Figure 1(e). Results obtained for estimates of posterior mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ with respect to the different prior combinations for both the parametric and semi-parametric model in Tables 24 and 25. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2,3$ and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 19 in Appendix B.
Table 24: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 1.329 0.345 (0.639, 1.991) 1.343 0.384 (0.544, 2.079)
$\beta_{2}$ 4.626 0.350 (3.908, 5.284) 4.596 0.377 (3.812, 5.311)
$\beta_{3}$ 1.844 0.344 (1.104, 2.503) 1.837 0.349 (1.158, 2.513)
$\beta_{4}$ 3.958 0.339 (3.286, 4.638) 3.938 0.376 (3.201, 4.655)
$\beta_{5}$ 2.622 0.346 (1.940, 3.320) 2.608 0.370 (1.883, 3.305)
$\beta_{6}$ 9.643 0.352 (8.909, 10.326) 9.625 0.422 (8.598, 10.330)
$\beta_{7}$ 2.147 0.334 (1.476, 2.825) 2.133 0.373 (1.356, 2.882)
$\beta_{8}$ 5.694 0.352 (5.069, 6.394) 5.686 0.386 (4.863, 6.398)
$\beta_{9}$ 2.140 0.347 (1.456, 2.866) 2.108 0.368 (1.353, 2.795)
$\beta_{10}$ 8.207 0.338 (7.550, 8.889) 8.185 0.407 (7.308, 8.920)
$\alpha$ 2.052 0.110 (1.846, 2.264) 2.057 0.136 (1.827, 2.346)
$\sigma_{\epsilon}^{2}$ 0.994 0.122 (0.784, 1.264) 0.990 0.135 (0.714, 1.253)
Table 25: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 5
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 1.309 0.352 (0.643, 2.029) 1.327 0.342 (0.666, 1.995)
$\beta_{2}$ 4.620 0.365 (3.928, 5.338) 4.629 0.353 (3.994, 5.359)
$\beta_{3}$ 1.793 0.346 (1.105, 2.522) 1.829 0.366 (1.166, 2.567)
$\beta_{4}$ 3.950 0.354 (3.303, 4.649) 3.963 0.344 (3.326, 4.633)
$\beta_{5}$ 2.583 0.367 (1.879, 3.311) 2.609 0.358 (1.916, 3.360)
$\beta_{6}$ 9.725 0.355 (9.032, 10.425) 9.690 0.347 (8.961, 10.353)
$\beta_{7}$ 2.116 0.347 (1.461, 2.777) 2.133 0.358 (1.475, 2.858)
$\beta_{8}$ 5.734 0.369 (5.012, 6.431) 5.722 0.344 (5.019, 6.361)
$\beta_{9}$ 2.082 0.362 (1.379, 2.839) 2.111 0.350 (1.446, 2.865)
$\beta_{10}$ 8.281 0.353 (7.548, 8.943) 8.279 0.347 (7.630, 8.983)
$\alpha$ 2.052 0.115 (1.808, 2.276) 2.048 0.113 (1.826, 2.267)
$\sigma_{\epsilon}^{2}$ 1.016 0.126 (0.804, 1.301) 1.011 0.123 (0.788, 1.267)
It is evident from Tables 24 and 25 that for the proposed semi-parametric and parametric method seems to overestimate the true error variance. On the other side, 95% credible interval contains true fixed value for both the methods. The posterior mean for semi-parametric model is slightly higher than parametric method under prior 5. Similar to the previous cases, we find that the estimates produced by both the methods are not significantly different under different prior combination as the number of observations increases. We estimate residual lifetime distribution with respect to prior 2.
Table 26: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $m=31$
Unit True value Parametric Semi-parametric
median 95% Predictive interval median 95% Predictive interval
1 3.284 7.356 (3.482, 14.451) 7.342 (3.209, 15.710)
2 1.035 1.359 (0.832, 1.873) 1.376 (0.881, 1.972)
3 3.842 4.853 (2.912, 8.097) 4.956 (2.955, 8.178)
4 1.196 1.757 (1.111, 2.413) 1.779 (1.107, 2.422)
5 2.702 3.206 (2.054, 4.544) 3.199 (1.974, 4.711)
6 0.014 0.145 (0.001, 0.313) 0.146 (0.002, 0.333)
7 3.703 4.165 (2.528, 6.506) 4.146 (2.361, 6.701)
8 0.701 0.925 (0.491, 1.268) 0.920 (0.525, 1.348)
9 2.756 4.126 (2.438, 6.355) 4.237 (2.675, 6.704)
10 0.043 0.335 (0.116, 0.602) 0.341 (0.105, 0.612)
Table 27: KS distance between true and approximated residual lifetime distribution for Bayesian parametric method and proposed Bayesian semi-parametric method
. Unit $m=11$ $m=31$ Parametric Semi-parametric Parametric Semi-parametric 1 0.719 0.707 0.877 0.867 2 0.376 0.412 0.259 0.247 3 0.469 0.471 0.302 0.279 4 0.173 0.199 0.273 0.302 5 0.595 0.595 0.305 0.321 6 0.092 0.105 0.117 0.128 7 0.403 0.464 0.592 0.571 8 0.412 0.391 0.191 0.192 9 0.319 0.312 0.345 0.375 10 0.121 0.121 0.108 0.131
Table 28: Prediction accuracy for Parametric and Semi-Parametric method
. Scenario $M_{1}$ $M_{2}$ Parametric Semi-parametric Parametric Semi-parametric RMSE MAE RMSE MAE RMSE MAE RMSE MAE $n=10,m=11$ 1.176 0.832 1.181 0.824 1.013 0.797 0.950 0.767 $n=10,m=31$ 1.433 0.946 1.447 0.957 1.411 0.949 1.425 0.960 $n=30,m=11$ 1.536 1.019 1.488 1.003 1.501 1.012 1.488 1.003 $n=30,m=31$ 0.916 0.842 0.923 0.842 0.904 0.838 0.902 0.834 $n=50,m=11$ 0.967 0.811 0.975 0.810 0.916 0.794 0.991 0.811 $n=50,m=31$ 0.716 0.751 0.801 0.794 0.710 0.747 0.807 0.795
The predictive intervals of residual lifetime for the individual units is given in Table 26. For most of the units, the produced predictive intervals for both parametric and semi-parametric methods contain the true value. Although for unit 1, predictive interval for parametric method fails to contain true residual lifetime whereas semi-parametric method does. For unit 10 both the method fails to contain the true value. It is also evident that both parametric and semi-parametric method produces predictive interval of very large length for the most slowly degrading units, which indicates to the fact that both the model is unable to predict occurrences of failures which will happen after a significantly long time. The histograms of the generated residual lifetimes for units 1-3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method are provided in Figure 11 and 12 respectively in the Appendix A of this article.
Kolmogorov-Smirnov (KS) distance obtained for $m=11,31$ and $n=10$ is given in Table 27. For most of the slowly degrading units for example unit 1, 3 value of KS statistic is less for semi-parametric method compared to parametric method for $m=31$. For most of the moderate degrading units for example unit 8, both parametric and semi-parametric methods produces almost similar KS value as $m$ increases. On the other side, for fastest degrading units value of KS distance is less for parametric method. It is evident from Table 28 that, as the number of observations per unit increases for $n=30,50$, both RMSE and MAE decreases. It is found that the produced errors seems to similar with respect to both model under $M_{1}$ and $M_{2}$ as $n$ and $m$ both increases. In each scenarios for case 3,4 and 5, the estimated median of residual lifetime is higher than true value which is also observed for case 1 and 2. The reason is similar the distribution functions for the individual residual lifetime contains terms which are average of many distribution functions and longer upper tail.
7 Data Analysis
Now we demonstrate analysis of fatigue crack size dataset given in Lu & Meeker[3]. The measurements are taken for 21 units. The crack lengths are measured at interval of $0.01$ million of cycles. It is considered that failure occurs when crack length exceeded the threshold value $D=1.6$ inches (see Lawless and Crowder [7], Lu and Meeker[3], Giorgio et al. [19]). For each unit the initial crack size was 0.9 inch and data collection was terminated at the first inspection after a unit’s crack size reached 1.6 inches or censored after 0.12 million cycles, whichever came first. Degradation paths of the units are presented in Figure 2. It may be noted from the plot that some of the units are degrading faster than other units, so it can be considered that the units are coming from heterogeneous population. For semi-parametric method we take truncation number $N$ as 21.
Figure 2: Development of crack sizes over time [3]
Table 29: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,21$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2.
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 5.665 0.879 (4.010, 7.456) 5.690 0.790 (4.166, 7.368)
$\beta_{2}$ 5.617 0.812 (3.944, 7.206) 5.617 0.794 (4.073, 7.120)
$\beta_{3}$ 5.749 0.775 (4.280, 7.348) 5.759 0.741 (4.429, 7.362)
$\beta_{4}$ 5.653 0.794 (4.193, 7.398) 5.670 0.755 (4.171, 7.181)
$\beta_{5}$ 5.679 0.788 (4.271, 7.409) 5.633 0.763 (4.140, 7.247)
$\beta_{6}$ 5.626 0.757 (4.162, 7.171) 5.644 0.719 (4.212, 7.037)
$\beta_{7}$ 5.614 0.784 (4.038, 7.094) 5.539 0.756 (4.048, 7.110)
$\beta_{8}$ 5.585 0.792 (3.916, 7.270) 5.530 0.714 (4.230, 7.023)
$\beta_{9}$ 5.573 0.766 (4.048, 6.988) 5.563 0.703 (4.153, 6.980)
$\beta_{10}$ 5.455 0.757 (4.064, 6.904) 5.487 0.707 (4.131, 6.881)
$\beta_{11}$ 5.444 0.777 (3.964, 6.926) 5.412 0.720 (3.947, 6.837)
$\beta_{12}$ 5.415 0.770 (3.986, 6.902) 5.380 0.734 (3.928, 6.971)
$\beta_{13}$ 5.235 0.741 (3.762, 6.755) 5.296 0.762 (3.859, 6.900)
$\beta_{14}$ 5.146 0.757 (3.514, 6.549) 5.185 0.711 (3.739, 6.579)
$\beta_{15}$ 5.278 0.775 (3.764, 6.762) 5.218 0.718 (3.800, 6.683)
$\beta_{16}$ 5.037 0.802 (3.368, 6.638) 5.063 0.752 (3.588, 6.545)
$\beta_{17}$ 5.052 0.762 (3.549, 6.603) 5.027 0.729 (3.505, 6.435)
$\beta_{18}$ 4.934 0.807 (3.249, 6.428) 5.014 0.766 (3.373, 6.416)
$\beta_{19}$ 4.895 0.798 (3.192, 6.411) 4.943 0.733 (3.379, 6.352)
$\beta_{20}$ 4.886 0.787 (3.292, 6.357) 4.880 0.740 (3.328, 6.271)
$\beta_{21}$ 4.837 0.812 (3.094, 6.359) 4.835 0.739 (3.316, 6.164)
$\alpha$ 0.845 0.031 (0.783, 0.905) 0.844 0.028 (0.787, 0.896)
$\sigma_{\epsilon}^{2}$ 0.114 0.019 (0.077, 0.152) 0.113 0.017 (0.082, 0.152)
The mean, standard deviation (s.d) and 95% credible interval (C.I) of posterior samples of random effects for each of 21 units, fixed effect $\alpha$ and variance of measurement error $\sigma_{\epsilon}^{2}$ obtained for both parametric and semi-parametric method under prior 2, are presented in Table 29. Observe that for some units, posterior mean of random effect is much higher than some other units. For example for unit 1, the estimated posterior mean for random effects is much higher than that of unit 20, which can motivate us to analyze this data using the proposed semi-parametric model. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2$ and 3 and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 20 in Appendix B to this article.
Table 30: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $m=31$
Unit True value Parametric Semi-parametric
median 95% Predictive interval median 95% Predictive interval
1 0.008 0.07 (0.0003,0.191) 0.068 (0.0003, 0.175)
2 0.010 0.063 (0.0001, 0.174) 0.061 (0.0007, 0.164)
3 0.001 0.056 (0.0001, 0.152) 0.059 (0.0002, 0.153)
4 0.003 0.056 (0.0002, 0.148) 0.056 (0.0002, 0.158)
5 0.003 0.057 (0.0008, 0.162) 0.061 (0.0002, 0.160)
6 0.006 0.058 (0.0001, 0.156) 0.055 (0.0001, 0.153)
7 0.006 0.057 (0.0001, 0.163) 0.055 (0.0001, 0.150)
8 0.009 0.057 (0.0001, 0.167) 0.058 (0.0001, 0.151)
9 0.003 0.057 (0.0002, 0.152) 0.055 (0.0001, 0.157)
10 0.005 0.051 (0.0001, 0.153) 0.059 (0.0002, 0.161)
11 0.008 0.059 (0.0003, 0.157) 0.056 (0.0001, 0.159)
12 0.008 0.056 (0.0001, 0.151) 0.057 (0.0001, 0.164)
13 0.009 0.063 (0.0001, 0.161) 0.055 (0.0001, 0.159)
14 0.013 0.056 (0.0001, 0.167) 0.055 (0.0006, 0.161)
15 0.018 0.057 (0.0004, 0.158) 0.054 (0.0003, 0.163)
16 0.024 0.062 (0.0002, 0.178) 0.053 (0.0001, 0.166)
17 0.026 0.060 (0.0002, 0.165) 0.063 (0.0007, 0.174)
18 0.031 0.062 (0.0001, 0.185) 0.062 (0.0001, 0.178)
19 0.040 0.068 (0.0002, 0.189) 0.063 (0.0002, 0.172)
20 0.047 0.068 (0.0002, 0.192) 0.067 (0.0001, 0.183)
21 0.050 0.069 (0.0003, 0.191) 0.066 (0.0001, 0.186)
We derive residual lifetime distribution of individual units corresponding to the last time point at which degrading unit has not crossed the threshold value with respect each prior combination. Table 30 presents residual lifetime estimates where the lower limit of the predictive interval is rounded off to four decimal places and rest of the values are rounded off to three decimal places along with the true values for each of 21 units. Histograms of the generated residual lifetimes of units 1-3 along with the trace plot and autocorrelation plot for the generated samples for semi-parametric and parametric method are provided in Figures 13 and 14 respectively in the Appendix A to this article. It is evident from Table 31 and 32 that both proposed semi-parametric and parametric method perform equally well in almost all the cases. Also, the choice of prior is not impacting the prediction accuracy for both models. It is also evident that for most of the cases the produced predictive intervals for both the methods contain true value for the residual lifetime.
Table 31: Prediction accuracy for Parametric method
Scenario $M_{1}$ $M_{2}$
RMSE MAE RMSE MAE
Prior 1 0.047 0.209 0.048 0.211
Prior 2 0.046 0.212 0.047 0.211
Prior 3 0.045 0.201 0.046 0.202
Table 32: Prediction accuracy for Semi-parametric method
Scenario $M_{1}$ $M_{2}$
RMSE MAE RMSE MAE
Prior 1 0.047 0.206 0.045 0.201
Prior 2 0.046 0.209 0.048 0.214
Prior 3 0.045 0.207 0.045 0.207
Prior 4 0.045 0.204 0.044 0.200
Prior 5 0.046 0.206 0.047 0.207
@@ 8 Conclusion
The objective of this article is to predict residual lifetime of units in a heterogeneous situation. To deal with heterogeneity, we proposed a Bayesian semi-parametric degradation model. In the first part we assumed a general path model where the random effects are modeled using the Dirichlet process mixture of normal distributions. Model hierarchy is represented according to Ishwaran & Zarrerpur [49]. Gibbs sampling is used to draw samples from the posterior distribution of the model parameters. In the second part of the problem we developed the residual lifetime distribution of each unit which depends on the samples drawn from posterior distribution. Finally we simulate samples from this distribution using MCMC technique and considered the sample median as the predicted residual lifetime of each unit. For implementation of the proposed method we use the R programming language. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method with respect to the error criteria discussed earlier in this paper. It is found that in both parametric and semi-parametric model, the estimated median of residual lifetime is higher than true value for most of the cases. The reason for this is the distribution functions derived to simulate observations involves terms which are average of many distribution functions and as a result this induces a much longer upper tail. In simulation study it is found that, in most of the cases the proposed method is performing well compared to the parametric method for the the units which degrade slowly compared to others, whereas the parametric method performs better than the proposed method for the fastest degrading units in most of scenarios. Finally, we applied the model to Fatigue Crack-Size dataset to evaluate the performance of our model. In this case, it is found that the prediction is almost similar for both the proposed semi-parametric and parametric method, but for most of the units, Bayesian semi-parametric method produces more accurate estimates for residual lifetime. The proposed Bayesian semi-parametric degradation model may be used for similar problems like in medical diagnosis contexts as well.
APPENDIX A : MCMC convergence for residual lifetime
Case 1
@@
Figure 3: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.
@@
Figure 4: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
Case 2
@@
Figure 5: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.
@@
Figure 6: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
Case 3
@@
Figure 7: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.
@@
Figure 8: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
Case 4
@@
Figure 9: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.
@@
Figure 10: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
Case 5
@@
Figure 11: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.
@@
Figure 12: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by parametric method.
Fatigue-Crack Size dataset
@@
Figure 13: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot and right side denotes autocorrelation plot of the samples produced for Unit 1-3 and $n=10$, $m=31$ by semi-parametric method.